OpenStudy (anonymous):

What is the range of the function y = -x^2 + 1?

OpenStudy (anonymous):

range and y are the same thing so find which number is the y

OpenStudy (anonymous):

y ≤ -1 y ≥ -1 I thnk this is the answer y ≤ 1 y ≥ 1

OpenStudy (anonymous):

well the range is 1 but I don't know which way the sign is faced.

OpenStudy (anonymous):

\[\ge or \le \]

OpenStudy (anonymous):

well is there a sign that is originally in the problem?

OpenStudy (anonymous):

i think it is \[\le \]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

hm then I am not sure because usually those signs arent thrown in aren't thrown in last second

OpenStudy (anonymous):

ya

Directrix (directrix):

y = -x^2 + 1 - (x) ^2 takes on a value of 0 or a negative number for any real value of x. So, the greatest value - (x)^2 can have is zero. Then, one is added to it. So, y = -x^2 + 1 maxes out when y = 1 and x is zero. The range is the set of all possible values of y. y cannot be bigger than 1. So, the range is ?